OpenStudy (anonymous):

Please Help! Show that the sum of the first n positive odd numbers is a square number.

geerky42 (geerky42):

Exactly how should you show it?

geerky42 (geerky42):

It's pretty easy to show it via picture (wordless proof)

OpenStudy (anonymous):

@geerky42 I wanted to know of how to show it. I was using a table but that did not help me so I started doing this 1+3=4 1+3+5=9 1+3+5+9=25 and so on

OpenStudy (anonymous):

you can certainly prove by induction

OpenStudy (anonymous):

induction?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what is that ?

OpenStudy (anonymous):

nevermind